An electron enters a magnetic field of flux density 7.6*10^-3 and an electric field of intensity 5.6*10^3 Vm-1 .

a. Find speed of the electron after passing through the plate

b. Calculate radius of curvature of path in magnetic field only




1
Expert's answer
2021-08-11T09:53:42-0400

(a) the speed of the electron

"v=\frac{E}{B}=\frac{5.6*10^3\:\rm V/m}{7.6*10^{-3}\:\rm T}=7.4*10^5\:\rm m/s"

(b) the radius of curvature of path

"r=\frac{mv}{qB}\\
=\frac{9.1*10^{-31}\:\rm kg*7.4*10^5\:\rm m/s}{1.6*10^{-19}\:\rm C*7.6*10^{-3}\:\rm T}=5.5*10^{-4}\:\rm m"